OpenStudy (anonymous):

what is the length of the missing side of a right triangle whose hyothenuse is 15 and whose third side is 9?

Directrix (directrix):

12

Directrix (directrix):

Use the Pythagorean Theorem: 9^2 + x^2 = 15^2 81 + x^2 = 225 x^2 = 144 x = 12